OpenStudy (anonymous):

Find the area of the region bounded by the tangent line to the curve f(x)=x^.5 at x=4, the curve, and the x-axis

OpenStudy (anonymous):

i really just need help finding the tangent line's equation

OpenStudy (anonymous):

The slope of the tangent line is f'(4)

OpenStudy (anonymous):

ok, so how do i take that and make an equation?

OpenStudy (anonymous):

f'(x) = .5x^(-.5) f'(4) = .5*4^(-.5) = .5 / 2 = .25

OpenStudy (anonymous):

the tangent has to go through the point (4, 2)

OpenStudy (anonymous):

2 = 4*.25 + b b = 2 - 1 = 1 the tangent line is y = .25x + 1

OpenStudy (anonymous):

how do you know that?

OpenStudy (anonymous):

because that's the value of f(4)

OpenStudy (anonymous):

that it has to go throug (4, 2), i mean

OpenStudy (anonymous):

the tangent line must go through the same point as the function

OpenStudy (anonymous):

ohhh ok. thanks so much i understand that now

OpenStudy (anonymous):

Integrate[.25 x + 1 - x^.5, {x, 0, 4}] = 2/3

OpenStudy (anonymous):

ok thank you :)

OpenStudy (anonymous):

wait wouldnt you have to set the eq's equal to each other because its not bounded by the y axis?